OpenStudy (anonymous):

Which quadratic equation has roots -2, 1/5 A. x^2+4x+4=0 B. 5x^2-9x-2=0 C.5x^2+9x-2=0 D.5x^2-11x+2=0 How do you find the roots of an equation?

OpenStudy (anonymous):

you find roots of an equation by factoring the given equation

OpenStudy (anonymous):

i still don't understand how they got 1/5 as a root. i have narrowed it down between c and b

OpenStudy (lgbasallote):

there is an easier way actually... since -2 and 1/5 are roots that means.. x = -2 x = 1/5 so... (x + 2)(x - 1/5) is your polynomial now expand it

OpenStudy (anonymous):

ok, when you factor either c or b, you are left with two brackets...if you set each bracket = to zero and solve for x, you willl get the roots of the equation

OpenStudy (lgbasallote):

after you solve for that by FOIL/distributive..it seems you have to multiply every term by 5 to match your options

OpenStudy (anonymous):

i ended up getting c.

OpenStudy (anonymous):

in the problem, they give you roots of -2, and 1/5 ...those numbers are also what you get when you solve for x... x= -2 and x = 1/5 so, now, that you know the roots, or what "x" is equal to, you could work backward.. (x+2)(5x-1) ...this factoring gives you roots of -2 & 1/5 and if you use the FOIL method to multiply (x+2)(5x-1), you will get "C" as the answer!

OpenStudy (anonymous):

Thank you guys for all your help!

OpenStudy (anonymous):

I truly am highly appreciative.

OpenStudy (anonymous):

youur welcome! I'm glad I could help you better understand the stated problem! :)